OpenStudy (anonymous):

Draw the graph y=3x^2+4x Please, describe in detail

OpenStudy (anonymous):

(x-k)^2 =c(y-a)

OpenStudy (lgbasallote):

y -intercept let x = 0 y = 3(0)^2 + 4(0) therefore y-intercept = (0,0) now solve for x-intercept set y to 0 0 = 3x^2 + 4x solve for x 0 = x(3x + 4) so x = 0 and x = -4/3 does that give you any hints?

OpenStudy (anonymous):

Yes, but where is the vertex of the parabola? In (0;0) ?

OpenStudy (lgbasallote):

well to get vrtex you have to change it to the form akash said

OpenStudy (anonymous):

Take it to the completed square form

OpenStudy (anonymous):

change your equation into this form.. (x-k)^2 =c(y-a)

OpenStudy (anonymous):

@sauravshakya C=0 ?

OpenStudy (anonymous):

3(x^2 +4/3 x) = y

OpenStudy (anonymous):

(x+2/3)^2 = y/3 + 4/9

OpenStudy (anonymous):

y=3x^2+4x =3(x^2+4/3 x) =3{x^2+2*x*2/3 + (2/3)^2 -(2/3)^2} =3(x+2/3)^2-4/3

OpenStudy (anonymous):

(x+2/3)^2 = 1/3( y + 4/3)

OpenStudy (anonymous):

Now, Hint the vertex of a parabloa in the form y=a(x-h)^2+k is (h,k)

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

@sauravshakya I need the other way, because this way is not for school

OpenStudy (anonymous):

U mean u want to use direct formula

OpenStudy (anonymous):

What is direct formula?

OpenStudy (anonymous):

If y=ax^2+bx+c then x-coordinate of the parabola is -b/2a

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yes, -2/3

OpenStudy (anonymous):

YEP, thats x-coordinate and now find y-coordinate by substituting x=-2/3 in y=3x^2+4x

OpenStudy (anonymous):

okay, i understood, Thank you!

OpenStudy (anonymous):

WELCOMX